# Hop-Decorate: An Automated Atomistic Workflow for Generating Defect Transport Data in Chemically Complex Materials

### v1

@@ II.2 Implementation Details
Under the hood, HopDec uses LAMMPS Thompson et al. (2022) as the MD driver, requiring users to specify LAMMPS-style input parameters. For the case studies presented herein, the Cu-Ni study employed an embedded-atom method (EAM) Daw and Baskes (1984) interatomic potential for Cu-Ni alloys Foiles (1985), while the (Fe,Ni)Cr2O4 study utilized a Buckingham–Coulomb potential plus a Morse potential form by Chartier et al. Chartier et al. (2013) previously validated for defect transport simulations in spinel structures Hatton and Uberuaga (2024, 2023); Chartier et al. (2008); Van Brutzel et al. (2019).
@@
In Cu-Ni, defect identification was done through the process outlined in Figure 1 with centrosymmetry parameters of centroN=12 (N nearest neighbors) and centroCutoff=1 (below which atoms are considered to be in bulk). The complexity of the FeNiCr2O4 structure is such that no unique centrosymmetry parameters could reliably identify a lattice defect. Therefore, HopDec was only invoked in Redecoration mode with a pre-defined transition of the vacancy on the A (tetrahedral) sublattice. Many defect identification methodologies exist outside of centrosymmetry such as reference structure based approaches and geometric/tessellation based methods which would likely have greater success with complex structures such as this, however, these are currently not implemented in HopDec.
@@ III.3 Discussion
The (Fe,Ni)Cr2O4 spinel case study highlights both the capabilities and the challenges of applying HopDec to more structurally complex oxides. In the normal spinel configuration, migration barriers showed limited spread, consistent with minimal 1NN chemical variability around migrating defects. In contrast, the random spinel exhibited significant variability in migration energies and complex defect pathways, including the formation of intermediate split-vacancy states. Importantly, attempts to correlate migration energetics with simple metrics (e.g., 1NN/2NN species counts) were unsuccessful in the inverse spinel, indicating the need for more sophisticated descriptors, such as local order parameters or graph-based environment representations.
@@ IV Conclusion
The Hop-Decorate code is available as an open-source Python package at https://github.com/PeteHatton/Hop-Decorate under the MIT license. The repository includes example scripts and configuration files used in the case studies presented in this work. Users can reproduce all major results, extend the workflow for new materials systems, and integrate HopDec into high-throughput simulation environments. Any future updates or enhancements to the codebase will also be maintained through this repository.

### v2

@@ II.1 Hop-Decorate (HopDec)
The goal is to quantify the mobility of a given type of defect in a complex alloy. Ultimately, this information would be integrated into mesoscale models where the defect-defect interactions, such as recombination and clusters, would be handled. Thus, HopDec provides critical kinetic and thermodynamic information for such models.
@@ II.2 Implementation Details
Under the hood, HopDec uses LAMMPS Thompson et al. (2022) as the MD driver, requiring users to specify LAMMPS-style input parameters. For the case studies presented herein, the Cu-Ni study employed an embedded-atom method (EAM) Daw and Baskes (1984) interatomic potential for Cu-Ni alloys Foiles (1985), while the (Fe,Ni)Cr2O4 study utilized a Buckingham–Coulomb potential plus a Morse potential form by Chartier et al. Chartier et al. (2013) previously validated for defect transport simulations in spinel structures Hatton and Uberuaga (2024, 2023); Chartier et al. (2008); Van Brutzel et al. (2019). This interatomic potential assigns fixed partial charges to each species: Fe/Ni: 1.2+, Cr: 1.8+, and O: 1.2- reflecting the ionic behavior of the underlying spinel. As a result, defects such as vacancies and antisites are effectively charged. While the model does not account for charge transfer effects, this is not expected to significantly impact the properties studied here. In contrast to systems like Fe3O4 where Fe2+/Fe3+ redox plays a major role Srivastava et al. (2023), charge transfer in (Fe,Ni)Cr2O4 is believed to be minimal. Moreover, fixed-charge models have previously captured cation disorder in similar systems Yamamoto et al. (2008); Chartier et al. (2008); Van Brutzel et al. (2019); Hatton and Uberuaga (2023), suggesting that charge transfer has a limited effect on the electrostatic landscape driving disorder Stevanović et al. (2010). Additionally, contributions to the total energy from net charges in the system will cancel out identically for E and will not contribute to forces for NEB calculations.
@@
HopDec is initialized from an atomic structure containing the defect of interest in isolation, so that its kinetics and thermodynamics can be probed precisely as a function of varying local environment. In Cu-Ni, the defect identification was done through the process outlined in Figure 1 with centrosymmetry parameters of centroN=12 (N nearest neighbors) and centroCutoff=1 (below which atoms are considered to be in bulk). The complexity of the FeNiCr2O4 structure is such that no unique centrosymmetry parameters could reliably identify a lattice defect. Therefore, HopDec was only invoked in Redecoration mode with a pre-defined transition of the vacancy on the A (tetrahedral) sublattice. Many defect identification methodologies exist outside of centrosymmetry such as reference structure based approaches and geometric/tessellation based methods which would likely have greater success with complex structures such as this, however, these are currently not implemented in HopDec.
@@ III.3 Discussion
The (Fe,Ni)Cr2O4 spinel case study highlights both the capabilities and the challenges of applying HopDec to more structurally complex oxides. In the normal spinel configuration, migration barriers showed limited spread, consistent with minimal 1NN chemical variability around migrating defects. In contrast, the random spinel exhibited significant variability in migration energies and complex defect pathways, including the formation of intermediate split-vacancy states. Importantly, attempts to correlate migration energetics with simple metrics (e.g., 1NN/2NN species counts) were unsuccessful in the inverse spinel, indicting the need for more sophisticated descriptors, such as local order parameters or graph-based environment representations.
While these case studies do not represent the full compositional complexity typically associated with chemically complex alloys (CCAs) or high entropy systems, the HopDec framework is inherently general. Its design places no restrictions on the number of atomic species, types of sublattices, or crystal structure, and can be readily applied to multi-principal element systems. In fact, the rapid combinatorial growth of distinct local environments with increasing chemical complexity only reinforces the need for automated, scalable workflows like HopDec. These simpler systems were selected to clearly demonstrate and validate the methodology, but future work will focus on more chemically complex systems to fully showcase HopDec’s capabilities in the context of true CCAs.
@@ IV Conclusion
The Hop-Decorate code is available as an open-source Python package at https://github.com/PeteHatton/Hop-Decorate under the BSD-3 license. The repository includes example scripts and configuration files used in the case studies presented in this work. Users can reproduce all major results, extend the workflow for new materials systems, and integrate HopDec into high-throughput simulation environments. Any future updates or enhancements to the codebase will also be maintained through this repository.
@@ Acknowledgements.
- Srivastava et al. (2023) S. Srivastava, B. P. Uberuaga, and M. Asta, The Journal of Physical Chemistry C 127, 17460 (2023).
- Yamamoto et al. (2008) T. Yamamoto, A. Chartier, K. Yasuda, C. Meis, K. Shiiyama, and S. Matsumura, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 266, 2676 (2008).
- Stevanović et al. (2010) V. Stevanović, M. d’Avezac, and A. Zunger, Physical review letters 105, 075501 (2010).